OpenStudy (anonymous):

Which inequality does not have the solution {x| x > 5}? a)−4x<−20 b)6x<30 c)1/4x>5/4 d)−3/5x<−3

OpenStudy (mathhelp346):

Turn the sign into and equal sign 14+x=3 Subtract 14 from both sides x=-11

OpenStudy (mathhelp346):

*an

OpenStudy (anonymous):

\[x \le -11\]

OpenStudy (anonymous):

thank you this helped me alot! Would you mind helping me with another one?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Which inequality does not have the solution {x| x > 5}? \[a) -4x < -20 \] \[b) 6x < 30\] \[c) 1/4x > 5/4 \] \[d) -3/5x < -3\]

OpenStudy (anonymous):

a. \[-4x < -20\]\[\frac{ -4x }{ -4 } > \frac{ -20 }{ -4 }\]\[x > 5\]b.\[6x < 30\]\[\frac{ 6x }{ 6 } < \frac{ 30 }{ 6 }\]\[x < 5\]c\[\frac{ x }{ 4 } > \frac{ 5 }{ 4 }\]\[\frac{ 4x }{ 4 }> \frac{ 4 \times 5 }{ 4 }\]\[x > 5\]d. \[\frac{ -3x }{ 5 } < -3\]\[\frac{ 5 \times -3 x }{ -3 \times 5 } > \frac{ 5 \times -3 }{ -3 }\]\[x > 5\]

OpenStudy (anonymous):

so is it a?

OpenStudy (anonymous):

Look for the one that doesn't have x > 5

OpenStudy (anonymous):

oh i thought it was suppose to have that lol

OpenStudy (anonymous):

you have helped me a lot thank you

OpenStudy (anonymous):

You're welcome